# Sparse Representations of Dynamical Networks:

### v4

@@ Appendix D Proofs of the main results
Remark that, since the realization given by (82) is irreducible, then we have that $\begin{bmatrix}\widetilde{E}-\lambda\widetilde{A}&\widetilde{F}-\lambda\widetilde{B}\end{bmatrix}$ and $\begin{bmatrix}\widetilde{E}^{\top}-\lambda\widetilde{A}^{\top}&\widetilde{C}^{\top}\end{bmatrix}^{\top}$ have full row and, respectively, column rank $\0\forall\lambda\in\mathbb{C}\backslash\{0\}$. It is straightforward to check that $\begin{bmatrix}\widetilde{E}&\widetilde{F}\end{bmatrix}=\footnotesize\begin{bmatrix}I_{n-p}&O&O&O\\ O&O&I_{p}&O\end{bmatrix}$ has full row rank, while $\begin{bmatrix}\widetilde{E}^{\top}&\widetilde{C}^{\top}\end{bmatrix}^{\top}=\footnotesize\begin{bmatrix}I_{n-p}&O&A_{12}^{\top}\\ O&O&I_{p}\end{bmatrix}^{\top}$ has full column rank. Therefore, we conclude that (85) is an irreducible Rosebrock-type realization and we proceed to define the associated system matrix $\widetilde{\mathcal{S}}(\lambda):=\small\begin{bmatrix}\widetilde{E}-\lambda\widetilde{A}&\widetilde{F}-\lambda\widetilde{B}\\ \widetilde{C}&\widetilde{D}\end{bmatrix}$ in order to investigate the latter’s zeros at $\lambda=0$. Notice that $\normalsize\widetilde{\mathcal{S}}(0)=\scriptsize\begin{bmatrix}I_{n-p}&O&O&O\\ O&O&I_{p}&O\\ A_{12}&I_{p}&-A_{11}+A_{12}K&B_{1}\end{bmatrix}$ has full row rank and, thus, it follows that (85) has no zeros at $\lambda=0$, which implies that $\begin{bmatrix}\lambda I_{p}-\mathbf{W}(\lambda)&\mathbf{V}(\lambda)\end{bmatrix}$ does not have any zeros at infinity.
@@
Proof of Theorem 4.6 Point $\mathbf{a})$ boils down to the displacement of poles at infinity, via the procedure from Theorem 3.1 of [20]. We will apply this result using only state-space-based arguments, to avoid using the pencil realizations from [20].
@@
(86) $\normalsize\mathbf{\Theta}(\lambda)\begin{bmatrix}\lambda I_{p}-\mathbf{W}(\lambda)&\mathbf{V}(\lambda)\end{bmatrix}=\begin{bmatrix}\lambda\mathbf{\Theta}(\lambda)&O\end{bmatrix}+\mathbf{\Theta}(\lambda)\begin{bmatrix}\mathbf{W}(\lambda)&\mathbf{V}(\lambda)\end{bmatrix}\footnotesize\begin{bmatrix}-I_{p}&O\\ O&I_{m}\end{bmatrix}\normalsize,$
@@
The identity given in (87g) is obtained by introducing $n-p$ uncontrollable modes in the realization from (87c). We denote $A_{K}:=KA_{11}-KA_{12}K+A_{21}-A_{22}K$ and, using (15) and (16), we write
(88) $\mathbf{\Theta}(\lambda)\footnotesize\begin{bmatrix}\mathbf{W}(\lambda)&\mathbf{V}(\lambda)\end{bmatrix}\begin{bmatrix}-I_{p}&O\\ O&I_{m}\end{bmatrix}=\left[\footnotesize\begin{array}[]{cc|cc}A_{x}-\lambda I_{p}&B_{x}A_{12}&B_{x}(A_{12}K-A_{11})&B_{x}B_{1}\\ O&A_{22}+KA_{12}-\lambda I_{n-p}&-A_{K}&KB_{1}+B_{2}\\ \hline\cr C_{x}&O&O&O\end{array}\right].\normalsize$
@@
(90) $\footnotesize\begin{array}[]{r}(\lambda I_{p}-A_{x})U^{-1}\begin{bmatrix}-\mathbf{M}(\lambda)&\mathbf{N}(\lambda)\end{bmatrix}\end{array}=$ $\ (\lambda I_{p}-A_{x})\begin{bmatrix}I_{p}&O\end{bmatrix}\times$
×$\times\footnotesize\begin{bmatrix}(\lambda I_{p}-A_{x})^{-1}&(\lambda I_{p}-A_{x})^{-1}A_{12}(\lambda I_{n-p}-A_{22}-KA_{12})^{-1}\\ O&(\lambda I_{n-p}-A_{22}-KA_{12})^{-1}\end{bmatrix}\times$
(91) ×$\times\footnotesize\begin{bmatrix}A_{11}-A_{12}K+A_{x}&B_{1}\\ KA_{11}-KA_{12}K+A_{21}-A_{22}K&KB_{1}+B_{2}\end{bmatrix}+\begin{bmatrix}A_{x}-\lambda I_{p}&O\end{bmatrix}.$
@@
(93) $\footnotesize\begin{array}[]{r}(\lambda I_{p}-A_{x})U^{-1}\end{array}\footnotesize\begin{bmatrix}-\mathbf{M}(\lambda)&\mathbf{N}(\lambda)\end{bmatrix}=$ $\footnotesize\begin{bmatrix}A_{x}-\lambda I_{p}&O\end{bmatrix}+\footnotesize\begin{bmatrix}I_{p}&A_{12}(\lambda I_{n-p}-A_{22}-KA_{12})^{-1}\end{bmatrix}\times$
×$\qquad\qquad\quad\,\times\footnotesize\begin{bmatrix}A_{11}-A_{12}K-A_{x}&B_{1}\\ A_{K}&KB_{1}+B_{2}\end{bmatrix}$
(96) $=$ $\footnotesize\begin{bmatrix}-\lambda I_{p}&O\end{bmatrix}+\left[\footnotesize\begin{array}[]{c|cc}A_{22}+KA_{12}-\lambda I_{n-p}&A_{K}&KB_{1}+B_{2}\\ \hline\cr A_{12}&A_{11}-A_{12}K&B_{1}\end{array}\right].$
@@
$\begin{bmatrix}\frac{1}{\lambda}\mathbf{W}^{\top}(\lambda)\\ \frac{1}{\lambda}\mathbf{V}^{\top}(\lambda)\end{bmatrix}^{\top}$$\zeta$$w$$\begin{bmatrix}I_{m}\\ \mathbf{G}(\lambda)\end{bmatrix}$$v$$y-\zeta$$u$$z-r$$\delta_{u}$$r$
@@
Proof of Theorem 4.15 The proof boils down to showing that $\mathbf{K}_{d}(\lambda)$ is a stabilizing controller for $\begin{bmatrix}I_{p}&\mathbf{G}^{\top}(\lambda)\end{bmatrix}^{\top}$, considering a standard feedback loop as in Fig. 3. This represents a sufficient condition for the closed-loop interconnection shown in Fig. 5, located at the top of this page, to be internally stable. The desired result follows by noticing that Fig. 5 is equivalent to Fig. 4. In order to show that $\mathbf{K}_{d}(\lambda)$ is a stabilizing controller for $\begin{bmatrix}I_{p}&\mathbf{G}^{\top}(\lambda)\end{bmatrix}^{\top}$, we point out (see Section 5.3 of [40]) that an equivalent condition for this is the fact that all the entries of $\mathbf{H}(\lambda)$ are stable, where
(97) $\mathbf{H}(\lambda):=\small\begin{bmatrix}I_{p}\\ I_{p}\\ \mathbf{G}(\lambda)\end{bmatrix}\left(I_{p}-\begin{bmatrix}\frac{1}{\lambda}\mathbf{W}(\lambda)&\frac{1}{\lambda}\mathbf{V}(\lambda)\end{bmatrix}\begin{bmatrix}I_{p}\\ \mathbf{G}(\lambda)\end{bmatrix}\right)^{-1}\begin{bmatrix}I_{p}&\frac{1}{\lambda}\mathbf{W}(\lambda)&\frac{1}{\lambda}\mathbf{V}(\lambda)\end{bmatrix}.\normalsize$
@@
Since $\mathbf{K}(\lambda)$ is a stabilizing controller for $\mathbf{G}(\lambda)$, we have by point $5$ of Lemma 5.10 in [40] that $(\mathbf{M}_{\mathbf{K}}(\lambda)\mathbf{M}^{\top}_{\mathbf{G}}(\lambda)-\mathbf{N}_{\mathbf{K}}(\lambda)\mathbf{N}^{\top}_{\mathbf{G}}(\lambda))^{-1}$ must be stable. Finally, by (98), we have the fact that $\mathbf{\Theta}(\lambda)=\mathbf{M}_{\mathbf{K}}(\lambda)(\lambda I_{p}-\mathbf{W}(\lambda))^{-1}$ and, by noticing that $\mathbf{M}_{\mathbf{K}}(\lambda)$, $\mathbf{W}(\lambda)$ along with $\mathbf{\Theta}(\lambda)$ are all stable TFMs, we get that $\lambda\mathbf{\Theta}(\lambda)=\mathbf{M}_{\mathbf{K}}(\lambda)+\mathbf{\Theta}(\lambda)\mathbf{W}(\lambda)$ is also a stable TFM. With this fact established, return to (97) and rewrite it as
@@
×$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\times\left(\lambda I_{p}-\mathbf{W}(\lambda)\right)^{-1}\begin{bmatrix}\lambda I_{p}&\mathbf{W}(\lambda)&\mathbf{V}(\lambda)\end{bmatrix}$
@@
(101) $\small\begin{bmatrix}A_{d}-sI_{n}&B_{d}\end{bmatrix}=\footnotesize\begin{bmatrix}I_{p}&O\\ K&I_{n-p}\end{bmatrix}\footnotesize\begin{bmatrix}-sI_{p}&A_{11}-sI_{p}&A_{12}&B_{1}\\ sK&A_{21}&A_{22}-sI_{n-p}&B_{2}\end{bmatrix}\tiny\begin{bmatrix}I_{p}&O&-I_{p}&O\\ O&O&\phantom{-}I_{p}&O\\ O&I_{n-p}&-K&O\\ O&O&\phantom{-}O&I_{m}\end{bmatrix}.\normalsize$
@@
Proof of Theorem 4.21 The proof largely follows the same reasoning as the one belonging to Theorem III.6 in [24] and can be broken down into its two constituent parts: the discrete-time case $\mathbf{I)}$ and the continuous-time one $\mathbf{II)}$, respectively.
$\mathbf{I)}$ Begin with the discrete-time case and recall from Remark 4.17 that $\mathbf{K}_{d}(z)$ is stable, when starting from a stable SRTR pair, and that each of its rows is also a stable TFM. Thus, the stabilizable and detectable realizations for each such row, represented by $e_{i}^{\top}\mathbf{K}_{d}(z)=\left[\footnotesize\begin{array}[]{c|c}\widetilde{A}_{di}-zI_{n_{i}}&\widetilde{B}_{di}\\ \hline\cr\vspace{-1.5mm}\hfil\\ \widetilde{C}_{di}&\widetilde{D}_{di}\end{array}\right]$, possess pole pencils $\widetilde{A}_{di}-zI_{n_{i}}$ which have all their eigenvalues located in $\mathbb{S}$. Using these $p$ representations, form the realization
@@
Due to its block-diagonal structure, the pole pencil of the realization from (102) will have all its eigenvalues located in $\mathbb{S}$, which ensures that the obtained realization for $\mathbf{K}_{d}(z)$ will be stabilizable and detectable, while Theorem 4.15 shows that $\mathbf{K}_{d}(z)$ is a stabilizing controller for $\begin{bmatrix}I&\mathbf{G}^{\top}(z)\end{bmatrix}^{\top}$. Recall that $\mathbf{G}(z)=\left[\footnotesize\begin{array}[]{c|c}A-zI_{n}&B\\ \hline\cr C&D\end{array}\right]$ has a stabilizable and detectable realization and let $\begin{bmatrix}I&\mathbf{G}^{\top}(z)\end{bmatrix}^{\top}=\left[\footnotesize\begin{array}[]{c|cc}A^{\top}-zI_{n}&O&C^{\top}\\ \hline\cr\vspace{-2mm}\hfil\\ B^{\top}&I_{m}&D^{\top}\end{array}\right]^{\top}$, with the latter having a realization possessing the same state variables as the realization of $\mathbf{G}(z)$ and which is both stabilizable and detectable (since so is the realization of $\mathbf{G}(z)$). For these realizations of $\mathbf{G}(z)$ and $\begin{bmatrix}I&\mathbf{G}^{\top}(z)\end{bmatrix}^{\top}$, it is straightforward to check that the closed-loop state dynamics in Figs. 4 and 5 are described by the same set of equations (with respect to the same sets of inputs and outputs). Thus, by the same arguments as those in part $\mathbf{(IV)}$ of the proof of Theorem III.6 in [24], we employ the discrete-time versions of Definition 5.2 and of Lemmas 5.2 and 5.3 from [40] for the closed-loop interconnection in Fig. 5 to get the desired result.
$\mathbf{II)}$ For the continuous-time case, start by recalling the realization from (100), which was shown in the proof of Proposition 4.19 to be minimal. Since the chosen SRTR pair is stable, Lemma C.1 ensures that $A_{22}+KA_{12}$ has all its eigenvalues located in $\mathbb{S}$. Thus, $\mathbf{K}_{d}(s)$ has $p$ unstable poles in $s=0$ along with $n-p$ poles located in $\mathbb{S}$. Obtain now a realization for each $e_{i}^{\top}\mathbf{K}_{d}(s)$ as follows
@@
We conclude that the realization from (103) is both stabilizable and detectable and, since $A_{22}+KA_{12}$ has all its eigenvalues located in $\mathbb{S}$, each $e_{i}^{\top}\mathbf{K}_{d}(s)$ has one pole in $s=0$. Therefore, the pole pencil of any stabilizable and detectable realization belonging to $e_{i}^{\top}\mathbf{K}_{d}(s)$ will have only one unstable eigenvalue, at $s=0$. Compute now just such realizations for each $e_{i}^{\top}\mathbf{K}_{d}(s)=\left[\footnotesize\begin{array}[]{c|c}\widetilde{A}_{di}-sI_{n_{i}}&\widetilde{B}_{di}\\ \hline\cr\vspace{-1.5mm}\hfil\\ \widetilde{C}_{di}&\widetilde{D}_{di}\end{array}\right]$ and use them to form
@@
The pole pencil of the realization from (104) will have, due to its block-diagonal structure, exactly $p$ eigenvalues at $s=0$, with the rest of them being located in $\mathbb{S}$. However, recall that $\mathbf{K}_{d}(s)$ has only $p$ unstable poles, with all of them being located at $s=0$. Thus, by part $\mathbf{(II)}$ in the proof of Theorem III.6 from [24], the obtained realization of $\mathbf{K}_{d}(s)$ must be both stabilizable and detectable and, once again, Theorem 4.15 shows that $\mathbf{K}_{d}(s)$ is a stabilizing controller for $\begin{bmatrix}I&\mathbf{G}^{\top}(s)\end{bmatrix}^{\top}$. Recall the fact that $\mathbf{G}(s)=\left[\footnotesize\begin{array}[]{c|c}A-sI_{n}&B\\ \hline\cr C&D\end{array}\right]$ has a stabilizable and detectable realization and let $\begin{bmatrix}I&\mathbf{G}^{\top}(s)\end{bmatrix}^{\top}=\left[\footnotesize\begin{array}[]{c|cc}A^{\top}-sI_{n}&O&C^{\top}\\ \hline\cr\vspace{-1mm}\hfil\\ B^{\top}&I_{m}&D^{\top}\end{array}\right]^{\top}$, with the latter having a realization possessing the same state variables as the realization of $\mathbf{G}(s)$ and which is both stabilizable and detectable (since so is the realization of $\mathbf{G}(s)$). As in part $\mathbf{I)}$, the closed-loop state dynamics in Figs. 4 and 5 are described by the same set of equations (with respect to the same sets of inputs and outputs). By the same arguments as those in part $\mathbf{(IV)}$ of the proof of Theorem III.6 in [24], we employ Definition 5.2 and Lemmas 5.2 and 5.3 from [40] for the closed-loop interconnection in Fig. 5 to get the desired result.
@@
(105) $e_{p,i}^{\top}\small\begin{bmatrix}\mathbf{W}(\lambda)&\mathbf{V}(\lambda)\end{bmatrix}=\left[\footnotesize\begin{array}[]{c|cc}A_{22}+KA_{12}-\lambda I_{n-p}&KA_{11}-KA_{12}K+A_{21}-A_{22}K&KB_{1}+B_{2}\\ \hline\cr\vspace{-2mm}\hfil\\ e_{p,i}^{\top}A_{12}&e_{p,i}^{\top}(A_{11}-A_{12}K)&e_{p,i}^{\top}B_{1}\\ \end{array}\right].\normalsize$
@@
$\widetilde{B}_{i1}:=$ $\begin{bmatrix}I_{n-p-n_{i}}&O\end{bmatrix}Q_{i}\begin{bmatrix}KA_{11}-KA_{12}K+A_{21}-A_{22}K&KB_{1}+B_{2}\end{bmatrix},\vspace{-1mm}$
@@
(106) $e_{p,i}^{\top}\begin{bmatrix}\mathbf{W}(\lambda)&\mathbf{V}(\lambda)\end{bmatrix}=\left[\footnotesize\begin{array}[]{cc|c}\widetilde{A}_{11}-\lambda I_{n-p-n_{i}}&\widetilde{A}_{12}&\widetilde{B}_{i1}\\ O&\widetilde{A}_{22}-\lambda I_{n_{i}}&\widetilde{B}_{i2}\\ \hline\cr O&\left\|e_{p,i}^{\top}A_{12}\right\|e_{n_{i},n_{i}}^{\top}&\widetilde{D}_{i}\end{array}\right]=\left[\footnotesize\begin{array}[]{c|c}\widetilde{A}_{22}-\lambda I_{n_{i}}&\widetilde{B}_{i2}\\ \hline\cr\left\|e_{p,i}^{\top}A_{12}\right\|e_{n_{i},n_{i}}^{\top}&\widetilde{D}_{i}\end{array}\right],\vspace{-1mm}$
@@
Moreover, notice that if conditions $i)$ and $ii)$ from the result’s statement hold, then

### v7

@@ Appendix D Proofs of the main results
[…]
@@
Proof of Theorem 4.6 Point $\mathbf{a})$ boils down to the displacement of poles at infinity, via the procedure from Theorem 3.1 of [SIMAX]. We will apply this result using only state-space-based arguments, to avoid using the pencil realizations from [SIMAX].
@@
(86) $\normalsize\mathbf{\Theta}(\lambda)\begin{bmatrix}\lambda I_{p}-\mathbf{W}(\lambda)&\hskip-2.84526pt\mathbf{V}(\lambda)\end{bmatrix}=\begin{bmatrix}\lambda\mathbf{\Theta}(\lambda)&\hskip-2.84526ptO\end{bmatrix}+\mathbf{\Theta}(\lambda)\begin{bmatrix}\mathbf{W}(\lambda)&\hskip-2.84526pt\mathbf{V}(\lambda)\end{bmatrix}\footnotesize\begin{bmatrix}-I_{p}&O\\ O&I_{m}\end{bmatrix}\normalsize$,
@@
The identity given in (87g) is obtained by introducing npn-p uncontrollable modes in the realization from (87c). We denote $A_{K}:=KA_{11}-KA_{12}K+A_{21}-A_{22}K$ and, using (15) and (16), we write
(88) $\mathbf{\Theta}(\lambda)\footnotesize\begin{bmatrix}\mathbf{W}(\lambda)&\hskip-5.69054pt\mathbf{V}(\lambda)\end{bmatrix}\begin{bmatrix}-I_{p}&O\\ O&I_{m}\end{bmatrix}\hskip-2.84526pt=\hskip-2.84526pt\left[\footnotesize\begin{array}[]{cc|cc}\hskip-5.69054ptA_{x}-\lambda I_{p}&\hskip-8.53581ptB_{x}A_{12}&\hskip-2.84526ptB_{x}(A_{12}K-A_{11})&\hskip-8.53581ptB_{x}B_{1}\hskip-5.69054pt\\ \hskip-5.69054ptO&\hskip-8.53581ptA_{22}+KA_{12}-\lambda I_{n-p}\hskip-2.84526pt&-A_{K}&\hskip-8.53581ptKB_{1}+B_{2}\hskip-5.69054pt\\ \hline\cr\hskip-5.69054ptC_{x}&\hskip-8.53581ptO&O&\hskip-8.53581ptO\hskip-5.69054pt\end{array}\right].\normalsize$
@@
(90) $\footnotesize\begin{array}[]{r}(\lambda I_{p}-A_{x})U^{-1}\hskip-2.84526pt\begin{bmatrix}-\mathbf{M}(\lambda)&\hskip-5.69054pt\mathbf{N}(\lambda)\end{bmatrix}\end{array}\hskip-5.69054pt\hskip-2.84526pt$= $\hskip-2.84526pt\ (\lambda I_{p}-A_{x})\begin{bmatrix}I_{p}&O\end{bmatrix}\times$
$\hskip-2.84526pt\times\hskip-2.84526pt\footnotesize\begin{bmatrix}(\lambda I_{p}-A_{x})^{-1}&\hskip-5.69054pt(\lambda I_{p}-A_{x})^{-1}A_{12}(\lambda I_{n-p}-A_{22}-KA_{12})^{-1}\\ O&(\lambda I_{n-p}-A_{22}-KA_{12})^{-1}\end{bmatrix}\hskip-2.84526pt\times$
(91) $\hskip-2.84526pt\times\hskip-2.84526pt\footnotesize\begin{bmatrix}A_{11}-A_{12}K+A_{x}&B_{1}\\ KA_{11}-KA_{12}K+A_{21}-A_{22}K&\hskip-2.84526ptKB_{1}+B_{2}\end{bmatrix}\hskip-2.84526pt+\hskip-2.84526pt\begin{bmatrix}A_{x}-\lambda I_{p}&\hskip-5.69054ptO\end{bmatrix}$.
@@
(93) $\footnotesize\begin{array}[]{r}(\lambda I_{p}-A_{x})U^{-1}\end{array}\hskip-5.69054pt\footnotesize\begin{bmatrix}-\mathbf{M}(\lambda)&\hskip-5.69054pt\mathbf{N}(\lambda)\end{bmatrix}$= $\footnotesize\begin{bmatrix}A_{x}-\lambda I_{p}\hskip-5.69054pt&O\end{bmatrix}+\footnotesize\begin{bmatrix}I_{p}&A_{12}(\lambda I_{n-p}-A_{22}-KA_{12})^{-1}\end{bmatrix}\times$
$\qquad\qquad\quad\,\times\footnotesize\begin{bmatrix}A_{11}-A_{12}K-A_{x}&B_{1}\\ A_{K}&KB_{1}+B_{2}\end{bmatrix}$
(96) $=$ $\footnotesize\begin{bmatrix}-\lambda I_{p}\hskip-5.69054pt&O\end{bmatrix}+\left[\footnotesize\begin{array}[]{c|cc}\hskip-5.69054ptA_{22}+KA_{12}-\lambda I_{n-p}&A_{K}&\hskip-5.69054ptKB_{1}+B_{2}\hskip-5.69054pt\\ \hline\cr A_{12}&A_{11}-A_{12}K&\hskip-5.69054ptB_{1}\end{array}\right]\hskip-2.84526pt.\hskip-2.84526pt$
@@
[…]
@@
Proof of Theorem 4.15 The proof boils down to showing that $\mathbf{K}_{d}(\lambda)$ is a stabilizing controller for $\begin{bmatrix}I_{p}&\mathbf{G}^{\top}(\lambda)\end{bmatrix}^{\top}$, considering a standard feedback loop as in Fig. 3. This represents a sufficient condition for the closed-loop interconnection shown in Fig. 5, located at the top of this page, to be internally stable. The desired result follows by noticing that Fig. 5 is equivalent to Fig. 4. In order to show that $\mathbf{K}_{d}(\lambda)$ is a stabilizing controller for $\begin{bmatrix}I_{p}&\mathbf{G}^{\top}(\lambda)\end{bmatrix}^{\top}$, we point out (see Section 5.3 of [zhou]) that an equivalent condition for this is the fact that all the entries of $\mathbf{H}(\lambda)$ are stable, where
[…]
@@
Since $\mathbf{K}(\lambda)$ is a stabilizing controller for $\mathbf{G}(\lambda)$, we have by point $5$ of Lemma 5.10 in [zhou] that $(\mathbf{M}_{\mathbf{K}}(\lambda)\mathbf{M}^{\top}_{\mathbf{G}}(\lambda)-\mathbf{N}_{\mathbf{K}}(\lambda)\mathbf{N}^{\top}_{\mathbf{G}}(\lambda))^{-1}$ must be stable. Finally, by (98), we have the fact that $\mathbf{\Theta}(\lambda)=\mathbf{M}_{\mathbf{K}}(\lambda)(\lambda I_{p}-\mathbf{W}(\lambda))^{-1}$ and, by noticing that $\mathbf{M}_{\mathbf{K}}(\lambda)$, $\mathbf{W}(\lambda)$ along with $\mathbf{\Theta}(\lambda)$ are all stable TFMs, we get that $\lambda\mathbf{\Theta}(\lambda)=\mathbf{M}_{\mathbf{K}}(\lambda)+\mathbf{\Theta}(\lambda)\mathbf{W}(\lambda)$ is also a stable TFM. With this fact established, return to (97) and rewrite it as
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\times\left(\lambda I_{p}-\mathbf{W}(\lambda)\right)^{-1}\begin{bmatrix}\lambda I_{p}&\mathbf{W}(\lambda)&\mathbf{V}(\lambda)\end{bmatrix}$
@@
[…]
@@
Proof of Theorem 4.21 The proof largely follows the same reasoning as the one belonging to Theorem III.6 in [NRF] and can be broken down into its two constituent parts: the discrete-time case $\mathbf{I)}$ and the continuous-time one $\mathbf{II)}$, respectively.
$\mathbf{I)}$ Begin with the discrete-time case and recall from Remark 4.17 that $\mathbf{K}_{d}(z)$ is stable, when starting from a stable SRTR pair, and that each of its rows is also a stable TFM. Thus, the stabilizable and detectable realizations for each such row, represented by $e_{i}^{\top}\mathbf{K}_{d}(z)=\left[\footnotesize\begin{array}[]{c|c}\widetilde{A}_{di}-zI_{n_{i}}&\widetilde{B}_{di}\\ \hline\cr\vskip-4.2679pt\\ \widetilde{C}_{di}&\widetilde{D}_{di}\end{array}\right]$, possess pole pencils $\widetilde{A}_{di}-zI_{n_{i}}$ which have all their eigenvalues located in $\mathbb{S}$. Using these $p$ representations, form the realization
@@
Due to its block-diagonal structure, the pole pencil of the realization from (102) will have all its eigenvalues located in $\mathbb{S}$, which ensures that the obtained realization for $\mathbf{K}_{d}(z)$ will be stabilizable and detectable, while Theorem 4.15 shows that $\mathbf{K}_{d}(z)$ is a stabilizing controller for $\begin{bmatrix}I&\mathbf{G}^{\top}(z)\end{bmatrix}^{\top}$. Recall that $\mathbf{G}(z)=\left[\footnotesize\begin{array}[]{c|c}A-zI_{n}&B\\ \hline\cr C&D\end{array}\right]$ has a stabilizable and detectable realization and let $\begin{bmatrix}I&\mathbf{G}^{\top}(z)\end{bmatrix}^{\top}=\left[\footnotesize\begin{array}[]{c|cc}A^{\top}-zI_{n}&O&C^{\top}\\ \hline\cr\vskip-5.69054pt\\ B^{\top}&I_{m}&D^{\top}\end{array}\right]^{\top}$, with the latter having a realization possessing the same state variables as the realization of $\mathbf{G}(z)$ and which is both stabilizable and detectable (since so is the realization of $\mathbf{G}(z)$). For these realizations of $\mathbf{G}(z)$ and $\begin{bmatrix}I&\mathbf{G}^{\top}(z)\end{bmatrix}^{\top}$, it is straightforward to check that the closed-loop state dynamics in Figs. 4 and 5 are described by the same set of equations (with respect to the same sets of inputs and outputs). Thus, by the same arguments as those in part $\mathbf{(IV)}$ of the proof of Theorem III.6 in [NRF], we employ the discrete-time versions of Definition 5.2 and of Lemmas 5.2 and 5.3 from [zhou] for the closed-loop interconnection in Fig. 5 to get the desired result.
$\mathbf{II)}$ For the continuous-time case, start by recalling the realization from (100), which was shown in the proof of Proposition 4.19 to be minimal. Since the chosen SRTR pair is stable, Lemma C.1 ensures that $A_{22}+KA_{12}$ has all its eigenvalues located in $\mathbb{S}$. Thus, $\mathbf{K}_{d}(s)$ has $p$ unstable poles in $s=0$ along with npn-p poles located in $\mathbb{S}$. Obtain now a realization for each $e_{i}^{\top}\mathbf{K}_{d}(s)$ as follows
@@
We conclude that the realization from (103) is both stabilizable and detectable and, since $A_{22}+KA_{12}$ has all its eigenvalues located in $\mathbb{S}$, each $e_{i}^{\top}\mathbf{K}_{d}(s)$ has one pole in $s=0$. Therefore, the pole pencil of any stabilizable and detectable realization belonging to $e_{i}^{\top}\mathbf{K}_{d}(s)$ will have only one unstable eigenvalue, at $s=0$. Compute now just such realizations for each $e_{i}^{\top}\mathbf{K}_{d}(s)=\left[\footnotesize\begin{array}[]{c|c}\widetilde{A}_{di}-sI_{n_{i}}&\widetilde{B}_{di}\\ \hline\cr\vskip-4.2679pt\\ \widetilde{C}_{di}&\widetilde{D}_{di}\end{array}\right]$ and use them to form
@@
The pole pencil of the realization from (104) will have, due to its block-diagonal structure, exactly $p$ eigenvalues at $s=0$, with the rest of them being located in $\mathbb{S}$. However, recall that $\mathbf{K}_{d}(s)$ has only $p$ unstable poles, with all of them being located at $s=0$. Thus, by part $\mathbf{(II)}$ in the proof of Theorem III.6 from [NRF], the obtained realization of $\mathbf{K}_{d}(s)$ must be both stabilizable and detectable and, once again, Theorem 4.15 shows that $\mathbf{K}_{d}(s)$ is a stabilizing controller for $\begin{bmatrix}I&\mathbf{G}^{\top}(s)\end{bmatrix}^{\top}$. Recall the fact that $\mathbf{G}(s)=\left[\footnotesize\begin{array}[]{c|c}A-sI_{n}&B\\ \hline\cr C&D\end{array}\right]$ has a stabilizable and detectable realization and let $\begin{bmatrix}I&\mathbf{G}^{\top}(s)\end{bmatrix}^{\top}=\left[\footnotesize\begin{array}[]{c|cc}A^{\top}-sI_{n}&O&C^{\top}\\ \hline\cr\vskip-2.84526pt\\ B^{\top}&I_{m}&D^{\top}\end{array}\right]^{\top}$, with the latter having a realization possessing the same state variables as the realization of $\mathbf{G}(s)$ and which is both stabilizable and detectable (since so is the realization of $\mathbf{G}(s)$). As in part $\mathbf{I)}$, the closed-loop state dynamics in Figs. 4 and 5 are described by the same set of equations (with respect to the same sets of inputs and outputs). By the same arguments as those in part $\mathbf{(IV)}$ of the proof of Theorem III.6 in [NRF], we employ Definition 5.2 and Lemmas 5.2 and 5.3 from [zhou] for the closed-loop interconnection in Fig. 5 to get the desired result.
@@
(105) $e_{p,i}^{\top}\small\begin{bmatrix}\mathbf{W}(\lambda)&\hskip-5.69054pt\mathbf{V}(\lambda)\end{bmatrix}=\left[\footnotesize\begin{array}[]{c|cc}A_{22}+KA_{12}-\lambda I_{n-p}&KA_{11}-KA_{12}K+A_{21}-A_{22}K&KB_{1}+B_{2}\\ \hline\cr\vskip-5.69054pt\\ e_{p,i}^{\top}A_{12}&e_{p,i}^{\top}(A_{11}-A_{12}K)&e_{p,i}^{\top}B_{1}\\ \end{array}\right].\normalsize$
@@
[…]
@@
(106) $e_{p,i}^{\top}\begin{bmatrix}\mathbf{W}(\lambda)&\hskip-5.69054pt\mathbf{V}(\lambda)\end{bmatrix}\hskip-2.84526pt=\hskip-2.84526pt\left[\footnotesize\begin{array}[]{cc|c}\hskip-2.84526pt\widetilde{A}_{11}-\lambda I_{n-p-n_{i}}&\widetilde{A}_{12}&\widetilde{B}_{i1}\\ O&\widetilde{A}_{22}-\lambda I_{n_{i}}&\widetilde{B}_{i2}\hskip-2.84526pt\\ \hline\cr O&\left\|e_{p,i}^{\top}A_{12}\right\|e_{n_{i},n_{i}}^{\top}&\widetilde{D}_{i}\end{array}\right]\hskip-2.84526pt=\hskip-2.84526pt\left[\footnotesize\begin{array}[]{c|c}\widetilde{A}_{22}-\lambda I_{n_{i}}&\widetilde{B}_{i2}\hskip-2.84526pt\\ \hline\cr\hskip-2.84526pt\left\|e_{p,i}^{\top}A_{12}\right\|e_{n_{i},n_{i}}^{\top}&\widetilde{D}_{i}\end{array}\right]\hskip-2.84526pt,\vskip-2.84526pt$
@@
Moreover, notice that if conditions $i)$ and ii)ii) from the result’s statement hold, then